# OpsEval: A Comprehensive Benchmark Suite for Evaluating Large Language Models’ Capability in IT Operations Domain

### v4

@@ Abstract
The past decades have witnessed the rapid development of Information Technology (IT) systems, such as cloud computing, 5G networks, and financial information systems. Ensuring the stability of these IT systems has become an important issue. Large language models (LLMs) that have exhibited remarkable capabilities in NLP-related tasks are showing great potential in AIOps, such as root cause analysis of failures, generation of operations and maintenance scripts, and summarizing of alert information. Unlike knowledge in general corpora, knowledge of Ops varies with the different IT systems, encompassing various private sub-domain knowledge, sensitive to prompt engineering due to various sub-domains, and containing numerous terminologies. As a consequence, existing NLP-related benchmarks such as C-Eval and MMLU can not guide the selection of suitable LLMs for Ops, and current metrics like BLEU and ROUGE can not adequately reflect the question-answering (QA) effectiveness in the Ops domain. Therefore, this paper proposes a comprehensive benchmark suite named OpsEval, including an Ops-oriented evaluation dataset, an evaluation benchmark for Ops, and a specially designed QA evaluation method for Ops. Our dataset contains 7,184 multiple-choice questions and 1,736 QA questions. We have carefully selected and released 20% of the dataset by CC-BY-NC-4.0 license [^1] Data page is available at https://github.com/NetManAIOps/OpsEval-Datasets. written by domain experts in various sub-domains to assist current researchers in preliminary evaluations of their LLMs tailored for Ops (OpsLLM). The remaining undisclosed 80% of the data is used to prevent test set leakage. We test over 21 latest LLMs under various settings such as self-consistency, chain-of-thought, and in-context learning, revealing findings when applying LLMs to Ops. We also propose an evaluation method for QA in Ops, which has a coefficient of 0.9175 with human experts and is improved by 0.2470 and 1.313 compared to BLEU and ROUGE, respectively. Over the past seven months, our dataset and leaderboard [^2] Leaderboard is available at https://opseval.cstcloud.cn/content/leaderboard. have been continuously updated. For reproducibility, the evaluation framework code is publicly available.
1 Introduction
The IT Operations (Ops) plays a crucial role in maintaining the efficient and stable operation of information systems such as cloud computing, 5G networks [^3] Strictly speaking, 5G belongs to the field of communications technology (CT), but given its broad association with the information technology (IT) sector, for the sake of generality, we refer to it as IT operations, abbreviated as Ops, throughout the remainder of this paper. and financial information systems. As the Internet continues to expand rapidly, the scale and complexity of systems are escalating, leading to the emergence of artificial intelligence-assisted operations as a novel trend. Termed “AIOps” by Gartner (Lerner, 2017), this technique utilizes artificial intelligence to address (but is not limited to) tasks such as anomaly detection, fault analysis, generation of alert summaries, performance optimization, and capacity planning.
@@ 1 Introduction
In recent years, large language models (LLMs) have witnessed significant advancements. The latest models, such as GPT-4o (OpenAI, 2024), GPT-4V (OpenAI, 2023b), Meta-Llama-3 (AI@Meta, 2024), and GLM-4 (Zeng et al., 2022), have demonstrated exceptional generalization and task-planning capabilities. As a result, these models have provided numerous opportunities to enhance downstream domain-specific applications. With its advanced summarizing, report analyzing, and ability to diagnose errors, LLM is well suited for Ops on tasks like question answering, information summarizing, and report analysis. Hereinafter, we refer to the LLM used for Ops as OpsLLM, regardless of whether they have been optimized specifically for Ops.
@@
While there are benchmarks for assessing general-purpose NLP-related capabilities, such as C-EVAL (Huang et al., 2023) and AGIEval (Zhong et al., 2023), there are also benchmarks for specific domains, like FinEval (Zhang et al., 2023) in the financial sector and CMB (Wang et al., 2023a) in the medical sector. However, no benchmark exists to evaluate the effectiveness of LLMs or OpsLLMs in Ops tasks. There is an urgent need for an Ops benchmark that informs us about the performance of current LLMs on Ops tasks. On the other hand, a good benchmark can significantly aid the optimization process of OpsLLMs tailored for the Ops domain. Nevertheless, due to the specialty of the Ops tasks, constructing an Ops benchmark presents the following challenges:
@@
  • Sub-domains. There are many different sub-domains, such as 5G communications, cloud computing, and bank transaction systems. Ops within each sub-domain typically requires a combination of various capabilities, such as network configuration or terminology explanation, which we refer to as “tasks”. Due to the large number of sub-domains and tasks in the Ops field and the lack of an authoritative and systematic taxonomy, classifying a large number of questions becomes a challenge.
@@
  • QA metric. Existing metrics like BLEU only consider the similarity of model output to the reference in natural language aspects, which does not always reflect true performance in Ops tasks. Some terms and expressions in Ops sub-domains have specific meanings that LLM cannot summarize. Designing an automatic evaluation method that assesses the accuracy of QA of Ops from an accurate semantic level is challenging.
@@
To address these issues, we propose OpsEval, a comprehensive benchmark suite for evaluating LLMs’ capability in the IT operations domain. First, to tackle the challenge of benchmark data mostly being private and not publicly shareable, we initiated a community around AIOps, which has attracted dozens of companies to participate. We have selected 10 representative sub-domains from the community, allowing continuous data contributions from community members. We then aggregate data under the same sub-domain to ensure robustness in evaluation. Additionally, we generated multi-choice (MC) and question-answering (QA) questions as supplements based on publicly available network management books. To address the challenge of classifying the numerous sub-domains and tasks in the Ops field, we employ model-based pre-clustering and manual review to annotate eight tasks and three abilities for independent evaluation. Considering the prompt sensitivity of benchmark results, we systematically test model performance under self-consistency (SC), chain-of-thought (CoT), and few-shot in-context learning (ICL). The prompts used in our evaluation are also disclosed in the paper. Lastly, to address the inaccuracy of existing metrics in Ops QA evaluation, we design FAE-Score, which assesses questions from the perspectives of fluency, accuracy, and evidence. Experimental results show that our designed metrics align strongly with the annotations of human experts.
The contributions of our paper are as follows: 1) We introduce OpsEval, the first bilingual multi-task dataset in the operations domain, covering 8 tasks and 3 abilities with 8920 questions. To assist researchers in preliminary evaluating their OpsLLMs, we have carefully selected and released 20% of QAs from our benchmark licensed under CC-BY-NC-4.0, with the remaining 80% of undisclosed data preventing unfair evaluations due to data leakage (Wei, et.al., 2023) 2) Based on the dataset, we introduce the OpsEval evaluation benchmark, conducting independent and robust evaluations with various prompting techniques and a specifically designed evaluation metric named FAE-Score. Compared to the commonly employed BLEU and ROUGE metrics, FAE-Score exhibits a more pronounced congruence with the evaluations of human experts. Specifically, FAE-Score attains a correlation coefficient 0.9175 with expert assessments, surpassing the coefficients of 0.6705 for BLEU and -0.3957 for ROUGE. 3) We released an online leaderboard that continuously updates mainstream LLMs’ performance on Ops tasks for the past 7 months. To ensure the reproducibility and reliability of this leaderboard, we made our evaluation framework’s code publicly available [^4] Evaluation framework code is available at https://github.com/NetManAIOps/OpenCompass-OpsQA.. 4) Based on the results of OpsEval evaluation, we provide key observations and practical lessons to help domain practitioners make decisions such as whether existing models are sufficiently applicable within a specific sub-domain, the necessity for fine-tuning and whether model quantization compromises the effectiveness.
2 Related Works
As LLMs evolve rapidly, their complex and varied capabilities are increasingly recognized. As a result, there is a growing trend towards evaluation benchmarks tailored specifically for LLMs. These can be divided into two categories: general ability benchmarks and domain-specific benchmarks.
Table 1: A comparison of OpsEval with other popular datasets/benchmarks.
MMLU HELM BIG-bench SEAL C-Eval AGIEval FLUE MultiMedQA CMB NetOps OpsEval
Ops Domain Dataset
Open-sourced Benchmark
Up-to-date Leaderboard
General ability benchmarks assess the general abilities of LLMs across various tasks. These tasks evaluate LLMs’ capacity for logical reasoning, general knowledge, common sense, and other similar abilities rather than being confined to a particular domain. MMLU (Hendrycks et al., 2021) is a benchmark designed to measure knowledge acquired during pretraining by evaluating models exclusively in zero-shot and few-shot settings, covering 57 subjects across STEM. HELM (Liang et al., 2022) employs seven distinct metrics in 42 unique scenarios, offering a comprehensive evaluation of LLMs’ capabilities across multiple dimensions. BIG-bench (Srivastava et al., 2022) comprises 204 tasks spanning a wide array of topics, with a particular focus on tasks deemed beyond the reach of current LLMs. SEAL (AI, 2024b) features private, expert evaluations of leading frontiers models. C-Eval (Huang et al., 2023) is a comprehensive Chinese evaluation suite designed to assess Chinese LLMs’ advanced knowledge and reasoning abilities rigorously. AGIEval (Zhong et al., 2023) curates authentic questions from examinations such as the Chinese College Entrance Exam (CCEE) and the SAT, constructing a fundamentally human-centric evaluation dataset.
Domain-specific benchmarks evaluate the abilities of LLMs to handle tasks in specific fields. These benchmarks require LLMs to possess specialized knowledge in a specific domain and to respond in a manner consistent with the cognitive patterns of that field. Despite the rapid progression of LLMs in specialized domains, the evaluation metrics for these specific areas have received less attention. FLUE (Shah et al., 2022) is an open-source comprehensive suite of benchmarks, including new benchmarks across 5 NLP tasks in financial domain. MultiMedQA (Singhal et al., 2022) is an extensive medical question-answering dataset, with questions derived from professional medical exams, research, and consultation records. CMB (Wang et al., 2023a) includes multi-choice questions (CMB-Exam) and complex clinical questions based on real case studies (CMB-Clin). NetOps (Miao et al., 2023) focuses on evaluations in the network field, which is relevant to the field of Ops. NetOps includes multi-choice questions in both English and Chinese and a few cloze tests and question-answering questions. However, they only focus on wired network operations and while the dataset is released, they lack a benchmark that continuously updates the leaderboard.
3 OpsEval Benchmark
Figure 1: The framework of OpsEval
@@ 3 OpsEval Benchmark
3.1 Data Collection
Our benchmark questions have been collected from various sources; we summarize them into four categories: company materials, certification exams, Ops textbooks, and automated generation. Each source is highly esteemed globally and reviewed by our Ops collaborators.
Company Materials. include production environment materials like Ops tickets and error logs , as well as internal documents and tests for Ops staff training. We have established cooperative relationships with 10 companies, covering various sectors like telecommunications, finance, and Ops service/tool providers, and received expert collaboration and Ops materials from them. The Appendix A.1 provides information about the companies and experts.
Automated Generation. To enhance the diversity and depth of our test set, we source QAs from authoritative books covering a range of Ops domains by extracting textbook contents and asking GPT4 to generate questions. In Appendix A.7, we discussed the methodology and challenges.
Table 2: The number of questions in OpsEval, grouped by their sub-domains.
Sub-domain Source License Type Questions
Wired Network Operation Textbooks Public Multi-Choice 3901
5G Communication Certification Exams Public Multi-Choice 2615
Question-Answering 1162
Oracle Database Company Materials CC-BY-NC-4.0 Multi-Choice 497
Log Analysis Company Materials CC-BY-NC-4.0 Question-Answering 420
DevOps Automated Generation CC-BY-NC-4.0 Question-Answering 154
Securities Information System Company Materials CC-BY-NC-4.0 Multi-Choice 91
Hybrid Cloud Company Materials CC-BY-NC-4.0 Multi-Choice 40
Financial IT Company Materials CC-BY-NC-4.0 Multi-Choice 40
3.2 Preprocessing
Dependance Filtering: We have filtered out questions that rely on external images or document content to ensure the completeness of the question content itself. The filtering process was done by two parallel lists of empirical keywords in the question stems and the responses of GPT-3.5-turbo. The keyword list can be found in the Appendix A.2.
Question Categorization: We devise a categorization that captures many tasks that professionals confront in practical applications. The categorization process consists of two steps: automated screening and manual review. We first use GPT-4 for topic modeling to gain rough insights about the dataset and determine the relevance of each question to Ops, which resulted in more than 20 tasks but had an imbalanced distribution [^5] For prompts used during the preprocessing, please see in Appendix.. We then involved dozens of experts during the manual review process to categorize the questions into eight tasks and three abilities. The distribution of the questions across these eight tasks and three ability levels is shown in Table 3, and the details of each task and ability can be found in Appendix A.3.
Table 3: The distribution of different tasks and abilities of questions in OpsEval.

### v5

@@ Abstract.
In recent decades, the field of software engineering has driven the rapid evolution of Information Technology (IT) systems, including advances in cloud computing, 5G networks, and financial information platforms. Ensuring the stability, reliability, and robustness of these complex IT systems has emerged as a critical challenge. Large language models (LLMs) that have exhibited remarkable capabilities in NLP-related tasks are showing great potential in AIOps, such as root cause analysis of failures, generation of operations and maintenance scripts, and summarizing of alert information. Unlike knowledge in general corpora, knowledge of Ops varies with the different IT systems, encompassing various private sub-domain knowledge, sensitive to prompt engineering due to various sub-domains, and containing numerous terminologies. Existing NLP-related benchmarks can not guide the selection of suitable LLMs for Ops (OpsLLM), and current metrics (e.g., BLEU, ROUGE) can not adequately reflect the question-answering (QA) effectiveness in the Ops domain. We propose a comprehensive benchmark suite, OpsEval, including an Ops-oriented evaluation dataset, an Ops evaluation benchmark, and a specially designed Ops QA evaluation method. Our dataset contains 7,334 multiple-choice questions and 1,736 QA questions. We have carefully selected and released 20% of the dataset written by domain experts in various sub-domains to assist current researchers in preliminary evaluations of OpsLLMs [^1] Data page is available at https://github.com/NetManAIOps/OpsEval-Datasets. We test over 24 latest LLMs under various settings such as self-consistency, chain-of-thought, and in-context learning, revealing findings when applying LLMs to Ops. We also propose an evaluation method for QA in Ops, which has a coefficient of 0.9185 with human experts and is improved by 0.4471 and 1.366 compared to BLEU and ROUGE, respectively. Over the past one year, our dataset and leaderboard have been continuously updated.
Large language models, Operations, Benchmark, Evaluation, Prompt engineering
journalyear: 2025copyright: ccconference: 33rd ACM International Conference on the Foundations of Software Engineering; June 23–28, 2025; Trondheim, Norwaybooktitle: 33rd ACM International Conference on the Foundations of Software Engineering (FSE Companion ’25), June 23–28, 2025, Trondheim, Norwaydoi: 10.1145/3696630.3728572isbn: 979-8-4007-1276-0/2025/06ccs: Software and its engineeringccs: Computing methodologies Artificial intelligence
1. Introduction
IT Operations (Ops) plays a pivotal role in ensuring efficient and reliable functioning of software systems, including cloud computing, 5G networks, and financial platforms. With the rapid expansion of the Internet, the scale and complexity of software systems have grown exponentially, making traditional operations increasingly challenging. To address these challenges, artificial intelligence-assisted operations have emerged as a transformative approach, often referred to as “AIOps” by Gartner (Lerner, 2017). AIOps uses artificial intelligence to tackle critical software engineering tasks such as anomaly detection, fault diagnosis, and performance optimization.
@@ 1. Introduction
In parallel, recent advances in large language models (LLMs) have further expanded the potential of intelligent solutions in software operations. The latest models, such as GPT-4o (OpenAI, 2024), GPT-4V (OpenAI, 2023b), Meta-Llama-3 (AI@Meta, 2024), and GLM-4 (Zeng et al., 2022), have demonstrated exceptional generalization and task planning capabilities. As a result, these models have provided numerous opportunities to enhance downstream domain-specific applications. With its advanced text generation ability, LLM is well suited for Ops on tasks like question answering, information summarizing, and report analysis. Hereinafter, we refer to the LLM used for Ops as OpsLLM, regardless of whether they have been optimized specifically for Ops.
@@
While there are benchmarks for assessing general-purpose NLP-related capabilities, no benchmark exists to evaluate the effectiveness of LLMs or OpsLLMs in Ops tasks. There is an urgent need for an Ops benchmark that informs us about the performance of current LLMs on Ops tasks. On the other hand, a good benchmark can significantly aid the optimization process of OpsLLMs tailored for the Ops domain. Nevertheless, due to the specialty of the Ops tasks, constructing an Ops benchmark presents the following challenges:
Figure 1. The framework of OpsEval.
[…]
[The framework of OpsEval.]The framework of OpsEval.
@@
  • Sub-domains. The Ops field spans many sub-domains, like 5G communications, cloud computing, and bank transactions, each requiring a mix of capabilities, or “tasks,” such as network configuration or terminology explanation. The sheer number of sub-domains and tasks, combined with the absence of a systematic taxonomy, makes classifying questions challenging.
@@
  • QA metric. Existing metrics like BLEU focus on linguistic similarity between model output and reference answers, which often fails to capture true performance in Ops tasks. In Ops, it’s essential to assess whether the model’s answers address key points in the reference and are supported by sufficient evidence, reflecting the precise meanings of domain-specific terms.
@@
To address these issues, we propose OpsEval, a comprehensive benchmark suite for evaluating LLMs’ capability in the IT Operations domain, focusing on tasks essential to maintaining and troubleshooting live systems in real-world production environments. First, to tackle the challenge of benchmark data mostly being private, we initiated an AIOps community, which has attracted dozens of companies to participate. We have selected 9 representative sub-domains from the community, allowing continuous data contributions from community members. We aggregate data under the same sub-domain to ensure robustness in evaluation. Additionally, we generated multi-choice (MC) and question-answering (QA) questions as supplements based on publicly available network management books. To address the challenge of classifying the numerous sub-domains and tasks in the Ops field, we employ model-based pre-clustering and manual review to annotate eight tasks and three abilities. Considering the prompt sensitivity of benchmark results, we systematically test model performance under self-consistency (SC), chain-of-thought (CoT), and few-shot in-context learning (ICL). Lastly, to address the inaccuracy of existing metrics in Ops QA evaluation, we design FAE-Score, which evaluates model responses based on fluency, accuracy, and evidence, with each criterion having its own dedicated assessment method.
The contributions of our paper are as follows:
  1. (1)
    We introduce OpsEval, the first bilingual multi-task dataset in the Ops domain, covering 8 tasks and 3 abilities with 9,070 questions. To assist researchers in preliminary evaluating their OpsLLMs, we have carefully selected and released 20% of QAs from our benchmark licensed under CC-BY-NC-4.0, with the remaining 80% of undisclosed data preventing unfair evaluations due to data leakage (Wei and et.al., 2023).
  2. (2)
    Based on the dataset, we introduce the OpsEval evaluation benchmark, conducting independent and robust evaluations with various prompting techniques and a specifically designed evaluation metric, FAE-Score. Compared to the commonly employed BLEU and ROUGE metrics, FAE-Score exhibits a more pronounced congruence with the evaluations of human experts. Specifically, FAE-Score attains a correlation coefficient 0.9175 with expert assessments, surpassing the coefficients of 0.6705 for BLEU and -0.3957 for ROUGE.
  3. (3)
    Based on the results of OpsEval evaluation, we provide key observations and practical lessons to help domain practitioners make decisions such as whether existing models are sufficiently applicable within a specific sub-domain, the necessity for fine-tuning and whether model quantization compromises the effectiveness.
2. OpsEval Benchmark
@@
2.1. Data Collection
Table 1. Overview of the question distribution in OpsEval by sub-domains, tasks and abilities.
(a) The number of questions in OpsEval, grouped by their sub-domains.
Sub-domain Source Type Questions
Wired Network Operation Textbooks MC 3901
5G Communication Certification Exams MC 2615
QA 1162
Oracle Database Company Materials MC 497
Log Analysis Company Materials QA 420
DevOps Company Materials QA 154
Private Cloud Company Materials QA 150
Securities Info. Company Materials MC 91
Hybrid Cloud Company Materials MC 40
Financial IT Company Materials MC 40
Total 9,070
(b) The distribution of different tasks and abilities of questions in OpsEval.